### v2

# The temperature and metallicity distributions of the ICM: insights with TNG-Cluster for XRISM-like observations

Abstract
The new era of high-resolution X-ray spectroscopy will significantly improve our understanding of the intra-cluster medium (ICM) by providing precise constraints on its underlying physical properties. However, spectral fitting requires reasonable assumptions on the thermal and chemical distributions of the gas. We use the output of TNG-Cluster, the newest addition to the IllustrisTNG suite of cosmological magnetohydrodynamical simulations, to provide theoretical expectations for the multi-phase nature of the ICM across hundreds of clusters (M) based upon a realistic model for galaxy formation and evolution. We create and analyse, in an observer-like manner, end-to-end XRISM/Resolve mock observations towards cluster centres. We then systematically compare the intrinsic temperature and Fe abundance of the simulated gas with the inferred ones from spectral fitting via a variety of commonly used spectral-emission models. Our analysis suggests that models with a distribution of temperatures, better describe the broad thermal distributions of the ICM, as predicted by TNG-Cluster, but still incur biases in the inferred temperature of keV (16th-84th percentiles). However, all spectral-emission models systematically underestimate the Fe abundance of the central ICM by 0.12 Solar ( 22 per cent), almost an order of magnitude higher than the abundance errors reported in the literature, primarily due to projection effects. Selecting only strong cool core clusters leads to minor improvements on inference quality, removing the majority of outliers but maintaining similar overall biases and cluster-to-cluster scatter.
keywords:
X-ray: galaxies: clusters – Galaxies: clusters – Galaxies: clusters: intracluster medium – methods: numerical1 Introduction
Galaxy clusters are excellent probes of the large-scale structure of the Universe due to their rarity, masses and preferential position at the knots of the cosmic web (e.g. Allen et al., 2011). Their extended gaseous atmospheres consisting of X-ray emitting plasma, known as the intra-cluster medium (ICM), play a key role in our understanding of the evolution of large structures, since processes such as mergers, gas accretion from the cosmic web and feedback all leave signatures in the distribution and properties of this gas.
Various X-ray telescopes (e.g. XMM-Newton, Chandra, Suzaku, SRG/eROSITA) have studied the ICM and its thermal and chemical properties across spatial and mass scales (e.g. Kravtsov & Borgani, 2012; Mernier et al., 2018; Walker et al., 2019; Mernier et al., 2023, and references therein). However, the low spectral resolution of the CCD detectors (120 eV) so far limits the scope of X-ray analyses. For example, at CCD resolution, the dynamical state of the gas is hard to assess, as it lacks the spectral resolution for a direct measurement of gas motions from the Doppler shifts of spectral lines. Similarly, neighbouring emission lines may get blended, complicating the recovery of e.g. the abundances of various elemental species and the extent of the multi-phase thermal structure of the gas. As a result, many simplifying assumptions, such as coupling individual abundances to Iron (Fe) or assuming single temperature gas, have been routinely made.
In the current and upcoming new era of high-resolution X-ray spectroscopy, from Hitomi to missions such as XRISM111https://www.xrism.jaxa.jp/en/ and newAthena, these issues can be addressed. For example, as shown for Perseus, Hitomi’s spectral resolution of 5 eV revealed a wealth of previously-unresolved spectral lines of different species in the keV band (Hitomi Collaboration et al., 2017). XRISM has further promised to break the degeneracies in the thermal structure of the gas by providing robust measurements of the various thermal components in clusters.
However, by overcoming many previous degeneracies, systematic uncertainties become a more prominent source of errors. Assumptions made during the analysis of the X-ray spectra – such as the assumed shape of the temperature and metallicity distributions of the ICM or on the atomic libraries describing the emission lines, as well as unavoidable projection effects – all impact inferences, and may lead to significant biases. This can be problematic, especially for science cases that are sensitive to the exact values of cluster properties, such as their metal budgets (e.g. Gastaldello et al., 2021, and references therein) and the X-ray observable-mass relations used for cosmological inference (e.g. most recently Ghirardini et al., 2024).
Unfortunately, the physical properties of the ICM are sufficiently complex that these issues cannot be assessed empirically. In this case, numerical simulations are useful tools. With the inclusion of ever more and more realistic physical processes, simulations provide an increasingly detailed view of the ICM (e.g. Nelson et al., 2024, and references therein). A number of previous works have quantified biases on the observationally inferred properties of clusters by mocking the X-ray emission of simulated clusters and by focusing on past, current and future X-ray missions (e.g. Gardini et al., 2004; Rasia et al., 2006, 2008; Biffi et al., 2018; Cucchetti et al., 2018; Roncarelli et al., 2018; Mernier et al., 2020; Barnes et al., 2021; ZuHone et al., 2023b; Castellani et al., 2024; Veronesi et al., 2024; Zhang et al., 2024; Sanders, 2023, and references therein). However, no analysis has yet quantified the systematic effects due to specific parametric model choices, such as the ICM temperature distribution, when fitting high-resolution X-ray spectra.
In this work, we aim to assess the ability of spectral-emission models that are routinely adopted in the field to infer the mass-weighted temperature and Fe abundance of the ICM in cluster centres. Specifically, we attempt to quantify the systematic errors that might be introduced in said inferences due to the assumed description of the ICM’s temperature distribution, by simultaneously accounting for cluster-to-cluster variations. To do so, we use TNG-Cluster (Nelson et al., 2024), a recent extension of the IllustrisTNG simulations,222www.tng-project.org which returns a rich and statistically-large sample of simulated galaxy clusters well suited to study the ICM. Based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a) applied to a suite of zoom-in cosmological magneto-hydrodynamical simulations, TNG-Cluster not only creates galaxies at that are consistent with observations, but also – and of relevance to this work – produces the ICM in clusters with overall characteristics that are consistent with observations (Nelson et al., 2024).
It is thanks to the combination of numerical resolution and the adopted physical modelling that the output of TNG-Cluster is reasonably, if not quantitatively, realistic. This includes a mix of strong, weak and non cool cores (Lehle et al., 2024, 2025), realistic levels of gas velocity dispersion (Truong et al., 2024) and kinematics (Ayromlou et al., 2024), a non-negligible amount of cool gas (Rohr et al., 2024; Staffehl et al., 2025), a diversity of radio relics (Lee et al., 2024), and rich X-ray morphologies similar to observations (Chadayammuri et al., 2024), including cold fronts, shocks, and cavities (Truong et al., 2024; Prunier et al., 2024). Additionally, the ICM returned by TNG-Cluster exhibits thermodynamical profiles across the 352 clusters – density, temperature, entropy, and pressure – that are in the ballpark of observational inferences (Lehle et al., 2024). The distribution of cooling radii, derived from spectral fitting on both TNG-Cluster mocks and real Chandra observations, are quantitatively consistent (Prunier et al., 2025). And the metallicity profiles appear compatible with available observational results taken at face value (Nelson et al., 2024). Overall, the realism of the ICM produced by TNG-Cluster further supports the credibility of the analysis we set out.
With TNG-Cluster, we can hence examine the performance of the most commonly used temperature and metallicity models to fit observed X-ray spectra, as we can retrieve the “true” ICM properties directly from the simulations. Importantly, we do so across a sample of a few hundred clusters at any given redshift, hence providing robust expectations across the diverse cluster population. To make our analysis timely, we focus on the types of observations currently being taken with XRISM, specifically of the central regions of nearby galaxy clusters with the soft X-ray spectrometer Resolve, for which we develop an end-to-end mock and analysis pipeline that we apply to the simulated clusters.
This paper is arranged as follows. In Section 2.1 we present our sample of simulated galaxy clusters from TNG-Cluster. In Section 2.2 we describe our end-to-end mocking procedure, used to create XRISM/Resolve like observations of our input clusters, and to analyse them. In Section 3.1 we first examine the intrinsic properties of the ICM, as predicted by TNG-Cluster. We then compare the inferences of various spectral-emission modelling approaches on the temperature distribution and Fe abundance in Sections 3.2 and 3.3, respectively. We discuss the reasons and implications of our findings and expand on various effects in Section 4, and summarize in Section 5.
2 Data and methods
2.1 The TNG-Cluster simulation suite
TNG-Cluster is a collection of 352 zoom-in magnetohydrodynamical simulations of galaxy clusters with halo masses of log i.e. log. These are drawn from a cubic parent box of about 1 comoving Gpc per side. It complements the volume-limited TNG300 simulation (Springel et al., 2018; Pillepich et al., 2018b; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) with the highest halo masses. In particular, a subset of halos with log at have been randomly chosen for re-simulation in TNG-Cluster in order to produce a uniform selection, and hence to compensate for the halo mass drop in TNG300, while simultaneously including all halos with log from the same parent box. The volumes for re-simulations have been chosen such that the high-resolution zoom-in region around each cluster has a radius typically larger than two times the halo virial radius. For a more detailed description of the halo selection and model choices we refer to Nelson et al. (2024). Here we summarize additional salient aspects.
TNG-Cluster was simulated with the AREPO code (Springel, 2010) at a resolution of 1.1 M⊙ in gas/stellar target mass, matching that of the TNG300 fiducial run. It is fully based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a). This includes a variety of processes that are needed to simulate galaxy formation and evolution, such as stellar formation and evolution, metal enrichment, radiative cooling/heating, seeding and growth of SMBH and both stellar and SMBH (thermal and kinetic) feedback. Specifically, the IllustrisTNG model produces stellar particles, each representing a single-age stellar population with an initial mass function described by Chabrier (2003). Such populations evolve and return metals to the surrounding inter-stellar medium via AGB stars, type Ia and type II supernovae over time (Vogelsberger et al., 2013; Wiersma et al., 2009) based on the stellar yields described in detail in Pillepich et al. (2018a). The simulation directly tracks nine elements (H, He, C, N, O, Ne, Mg, Si, and Fe), in addition to Europium, tagging their production channel.
CDM cosmological parameters from the Planck observations of the cosmic microwave background (Planck Collaboration et al., 2016) are assumed in TNG-Cluster and throughout this paper: =0.3089, =0.6911, =0.0486, H0=67.74 km s-1 Mpc -1, =0.8159 and ns=0.9667, consistent with the other IllustrisTNG runs.
2.2 End-to-end XRISM/Resolve-like mocks of simulated ICM
In order to quantify the accuracy of the most commonly used X-ray emission models, we generate mock observations of our simulated clusters. In the following sections, we hence describe our end-to-end pipeline for XRISM/Resolve-like X-ray mock spectra using TNG-Cluster halos.
2.2.1 Modeling the intrinsic X-ray signal
First, we model the X-ray emission of the ICM based on the thermodynamical and chemical properties of the simulation gas cells, as predicted by TNG-Cluster. For the purposes of this work, we use the 4.4.0 version of the PyXSIM package (ZuHone et al., 2015), which follows the approach of the PHOX code (Biffi et al., 2012, 2013), in combination to the SOXS package (ZuHone et al., 2023a).
Similar to recent works (e.g. ZuHone et al., 2023b; Truong et al., 2024), for every cluster, we consider all the gas cells that are within a 3D radius of , including contribution from satellites, as emitting sources. While this approach excludes emission from structures outside that radius, ZuHone et al. (2023b) has shown for eROSITA mocks that their effect on the inferred cluster properties is negligible. Throughout, the centre of a cluster from TNG-Cluster is defined as the minimum of the potential. Furthermore, we assume that all gas cells satisfying the following criteria contribute to the X-ray emission of a given targeted cluster: have a temperature higher than K ( 0.03 keV), a negative net cooling rate, and are not star forming.
For each gas cell, we assume an optically-thin plasma in collisional ionisation equilibrium. We therefore model its emitted spectrum with an apec model (i.e. equivalently the vapec variant among XSPEC’s models) based on the AtomDB atomic library (v3.0.9 Smith et al., 2001; Foster et al., 2012).
The properties of the gas cells predicted within the TNG-Cluster simulation and relevant for this pipeline are: density, temperature, and absolute elemental abundances (w.r.t. Hydrogen). More specifically, for the latter, we use the TNG-Cluster outcome for all nine elements that the simulation tracks and a relative abundance ratio of for the rest (adopting Asplund et al., 2009), assuming that the absolute Fe abundance is a reasonable proxy of the underlying metal enrichment and that solar abundance ratios hold for those elements. Additionally, we account for the 3D position and velocity of each gas cell, to properly account for their effects to the properties of the generated X-ray photons. Namely, effects such as thermal broadening and Doppler shift of the emission lines are explicitly modelled by PyXSIM, producing an accurate representation of the X-ray emission from the ICM of our simulated galaxy clusters.
2.2.2 Observational strategy
In order to generate XRISM/Resolve mock event files, namely a list of detected photons of a given energy, from the simulated photons, we first project our halos along the z direction of the simulation box, representing a random direction for each cluster, and place them at the redshift of the Perseus Cluster (), as in Truong et al. (2024). Photoelectric absorption is also applied to the simulated photons, approximating with a wabs model the galactic absorption in the position of the Perseus cluster in the sky with a column density of (Kalberla et al., 2005).
Every cluster is observed with a single pointing aimed at each cluster’s centre and hence detecting the emission from all gas cells in a box volume: this hence represents a subset of the original ( 4-5 pMpc) re-simulation volume of the zoom simulations and is meant to replicate the XRISM/Resolve’s Field of View (FoV). In order to best approximate realistic observations of our targets, we also apply XRISM/Resolve’s instrumental responses to the intrinsic incoming photons. Namely, we use the Cycle 1 response files, as they have been incorporated in SOXS, for a high resolution ( 5eV) response matrix file (RMF) and a standard point-source ancillary response file (ARF) with the gate valve closed.
Regarding the background, as no instrumental background file is publicly available, we ignore it. On the other hand, for the foreground/background signals from the sky, we estimate that resolved point sources, the Cosmic X-ray Background (CXB), thermal emission from the Milky Way (MW) and the Local Hot Bubble (LHB) all provide a statistically-insignificant amount of detected photons and are effectively ignored in our mocks. This is a direct result of the small detector FoV, and its low efficiency in detecting low-energy photons due to the closed gate valve – their contribution is more relevant in the lower energy regime.
As a result of the procedure above, we obtain mock event files of XRISM/Resolve-like observations centred at the cores, unless otherwise specified, of the 352 clusters at from TNG-Cluster, assuming a reasonable exposure time of 100ks for each system.
Finally, we extract spectra from our event files using the dmcopy command from the Chandra Interactive Analysis of Observations (CIAO v.4.16) assuming a default spectral extraction, for the entirety of the FoV of the mock observation. A schematic representation of the result of our mock-observation pipeline can be seen in Figure 1.
2.3 Analysis i.e. fits of XRISM/Resolve-like spectra
For the purposes of our analysis, we will be using the X-ray spectral-fitting program XSPEC (Arnaud, 1996), using the AtomDB atomic database (v3.0.9), consistent with our X-ray mocks.
Throughout this work, we utilize four commonly-used spectral-emission models with different assumptions on the thermal structure i.e. temperature distribution of gas. The models in question are:
bvapec, as the baseline model commonly used to describe the gas: this assumes that the observed ICM is described by a single temperature, i.e. it returns a single-point estimate for its temperature;
double bvapec: this consists of two bvapec models and is primarily used to describe two distinct phases of gas (hot and cooler) assuming de facto a bimodal temperature distribution of the ICM;
bvgadem, which assumes a Normal distribution of ICM temperatures around a central peak;
bvlognorm333https://github.com/jeremysanders/lognorm, which assumes a Log-Normal distribution of ICM temperatures around a central peak.
For the Fe abundance, all adopted models assume that the observed ICM is described by a single value, as is the case for temperature with the bvapec and double bvapec models. We note that for all cases, both the line broadening due to the kinematics of the gas and the abundances of Si, S, Ar, Ca and Ni are taken into consideration, are modelled independently, and are simultaneously fit for. This is motivated by the fact that, if left fixed at a specific values, degeneracies between parameters could influence the model inferences on the temperature and Fe abundance. Consistently with Truong et al. (2024) and Xrism Collaboration et al. (2025c), our inferred velocity dispersions are comparable to the values obtained by central XRISM pointing observations (see Figure 15). However, further analysis and their exploitation are part of a future work. We note that H/He is kept fixed at 1 Solar and the abundances of all remaining elements are coupled to Fe, since they don’t have emission lines in our energy range.
For the purposes of this analysis, we do not deproject our spectra. With this choice we remain consistent with the spectral analysis of most recent XRISM observations (e.g. Xrism Collaboration et al., 2025a, b; Rose et al., 2025; Fujita et al., 2025). The only exception is that of XRISM Collaboration et al. (2025) for the Centaurus cluster, where a deprojected profile has been used for the interpretation of the results and has been constructed by fitting previous Chandra observations of Centaurus. An exact replication of this particular approach and study would require the full mocking and spectral fitting of Chandra-like observations for our entire sample, which would be beyond the scope of this project. Therefore, we choose to not account for projection effects on our spectral fitting, but we expand more on the implications of this choice in Section 4.2.1.
Regarding our spectral fitting strategy, all our extracted spectra are fitted in the [1.8-10] keV range. Even though the target redshift is known, we allow for the fit to vary it over a 50 per cent range in order to account for possible bulk motions of the gas along our line of sight. We use a guided approach to fit the model parameters gradually. However, it is still possible for the model to produce a non-valid fit. For this reason we introduce convergence criteria.
2.3.1 Acceptance criteria for spectral best fits
We distinguish between cases where the X-ray spectral fitting is successful (“accepted”) or not (”rejected”). The criteria are as follows.
For both the bvgadem and bvlognorm models, we require that the uncertainties on the central temperature and on the width of the distributions are smaller than the best-fit values themselves, as a minimum requirement for convergence. For the bvgadem model, we also require that the width of the assumed distribution is smaller than the best-fit central temperature, as otherwise the model would either imply the presence of negative temperatures or, due to its hard limit at 0, significantly alter its shape, in order to describe the thermal state of the gas. We note that this criterion is not as significant as the convergence of individual best-fit parameters, with 1 per cent of our clusters being rejected due to it. For the double bvapec model, we apply this same convergence criteria to each of the two temperature components, in order for the model to be meaningfully different from the single bvapec one. Lastly, we also apply a goodness-of-fit criterion to our spectra, requiring the C-statistic (Cash, 1979) of our spectral fits to deviate less than 3 from its expected Cash-statistic C value, following Kaastra (2017). If a best-fit result fails any of the above criteria, the fit is considered rejected; if it fulfils all of them, it is accepted.
3 Results
In the following sections we compare the ability of different spectral-fitting approaches to return the actual properties of the ICM, as predicted by TNG Cluster. We proceed with a multi-tiered approach.
Firstly, we compare the output temperature and Fe abundance of the best-fit spectral-emission models to the mass-weighted temperature and Fe abundance of the gas in the simulation located in a spherical volume centred at the centre of the cluster, with a radius of 1.5’ (33.5 kpc at the redshift of Perseus, to approximate XRISM FoV): we refer to this as gas in the central sphere or in the 3D cluster centre. Our choice is motivated by the fact that, observationally, one would hope that the output of the extracted spectrum represents the local conditions of the de-projected gas in the region of interest – for us, the XRISM FoV at the cluster centre.
Secondly, we compare the spectral-emission model inferences to the emission-weighted properties of all the gas in the adopted FoV for a central pointing, namely also considering all the foreground and background ICM gas: we refer to this as gas in a central pointing with a XRISM-like FoV. Previous studies have shown that the observed ICM abundance profiles can be better reproduced by weighting the properties of gas particles/cells with emission rather than mass (Biffi et al., 2017, 2018). Therefore, this can be considered a more apt comparison with our observables and a characterization of the “true” ICM state that is by construction closer to what the spectral fitting can return.
For the purposes of this analysis, we use the emission measure (EM) of each gas cell as its weight:
where , are the electron and hydrogen number densities and the volume of the gas cell, tracked by the simulation. We note that this weighting scheme differs from emission-weighted schemes that use the emissivity of each gas cell as their weight. Our choice in the weighting scheme is motivated by the fact that EM is indicative of each gas cell’s contribution to the spectrum, as the basis of the spectral-emission model’s normalisation. Furthermore, Cucchetti et al. (2018) have indicated that the metallicity of the gas, weighted by its EM, can be accurately recovered by fitting mock, high-resolution spectra.
3.1 The complexity of the thermal and chemical structure of cluster centres as predicted by TNG-Cluster
We first quantify the properties of the central X-ray emitting gas, in the ICM of our 352 targets from the TNG-Cluster simulation.
Temperature. Starting with the temperature distribution of the gas in the centre of our clusters (Figure 2, top), we find that, according to TNG-Cluster, the majority of the ICM mass is concentrated around a single temperature, the virial temperature of the cluster. The location of the peak exhibits, as expected, a strong correlation with halo mass, with the central ICM of more massive clusters being, on average, hotter. The width of the temperature distributions varies significantly across clusters, even between systems with similar halo masses. A subset of systems, independent of halo mass, appear to have extended tails of gas, extending to sub-keV temperatures. Considering that we are looking at the centrally-located gas, it is very likely that those tails are associated with the cooler inter-stellar (ISM) or circum-galactic (CGM) media of the brightest cluster galaxy (BCG; see also Rohr et al., 2024). A few clusters also exhibit very high, super-virial temperature tails, which, given the SMBH feedback model implemented in TNG-Cluster, may be driven by central energy injections (Truong et al., 2024). Overall, by visual inspection of the temperature distribution of each individual simulated cluster, we identify that at least 20 per cent of them (i.e. about 70 clusters of the 352) exhibit ICM in the central regions extending across more than one order of magnitude in temperature (in units of keV), i.e. with extended tails either towards cooler or super-virial temperatures.
Both the width and shape of the temperature distribution are relevant aspects for our analysis. So, what is the functional form of the temperature of the ICM at the centres of clusters favoured by TNG-Cluster?
We fit the temperature distributions in the 0.1-50 keV range with both the Normal and Log-Normal distributions. In Figure 3, we compare the geometric mean and best-fit temperature distributions of clusters in halo mass bins, for the intrinsic, mass-weighted gas in the 3D cluster centres (top) and for the emission-weighted gas in a XRISM-like FoV (bottom). We note that the average curves are not representative of each individual cluster’s best fit, but rather are meant to highlight possible systematic trends with cluster mass. The corresponding best-fit values of the Log-Normal distribution for each cluster are shown in Figure 4. The quality of each fit on a cluster-by-cluster basis is provided by the residuals in Figure 16, and it is summarized by the median across the entire cluster sample annotated as text in Figure 3.
As Figure 3 suggests, the temperature distribution cannot be perfectly described by either a Normal or a Log-Normal distribution in all systems. Both functional forms appear to be a reasonably good descriptor of the bulk of the temperature distribution around the virial temperature peaks. However, both models introduce biases primarily due to the presence of extended tails of gas in either direction, i.e. in those clusters where such tails are present. Both for the mass-weighted, centrally located gas as well as for the emission-weighted gas along the line of sight, the best-fit Normal distribution may underestimate to some small degree the amount of virial and hotter gas while typically systematically overestimating the amount of cooler gas, compared to the mode, present in the ICM.444Whereas the biases in the fits appear more pronounced in the most massive halos, this is an artifact of the averaging across multiple single-cluster distributions.
On the other hand, the Log-Normal model appears to be totally insensitive to the extended tails of sub-keV gas, which it fails to capture. This effect is also reflected in the , where, for both gas selections, the model tends to under-fit the true distribution of temperatures. Amid this overall picture, the cluster-to-cluster variation remains significant in all cases, as can be appreciated in Figure 16. From the visual inspection of the fits to each cluster temperature distribution, the Log-Normal model also tends to grossly underestimate the amount of gas at the hottest-end of the temperature distribution (in about 100 clusters of 352).
Overall, our analysis suggests that, a priori, neither functional form provides a perfectly accurate representation of the underlying thermal properties of the gas for all clusters. While both the Normal and Log-normal distributions are certainly more informative than a single or double delta function, they can not fully capture the tails towards cooler and hotter gas present in the intrinsic temperature distributions of a non-negligible fraction of clusters. It is evident that more sophisticated models are needed to fully capture the thermal structure of the ICM, but a more in depth analysis and the development of such models is outside the scope of this paper.
In Figure 4 we provide the best-fit central temperature and standard deviation for the Log-Normal fits, colour coded by the total mass of the cluster. Whereas there is a clear, and known, trend between average central ICM temperature and total cluster mass (spanning from a few keV to more than 10 keV for the studied clusters with log at ), the ICM in each cluster exhibits a diverse range of temperatures, irrespective of total mass. According to TNG-Cluster, the Log-Normal width of the central ICM temperature distribution varies between 0.4 and 10 keV.
Fe abundance. Regarding the Fe abundance of the ICM in the centres of our simulated clusters (Figure 2, bottom), TNG-Cluster predicts large system-to-system variations and broad distributions within individual clusters, across more than one order of magnitude in Solar units.
Upon inspection of all individual systems, the distributions can be broadly described as uni-modal, with the majority of the mass centred around a central peak for the largest majority of the clusters. In fact, we confirm that only a handful of metal-rich clusters exhibit secondary peaks at higher Fe abundances. However, unlike the temperature distributions, the peaks of the Fe abundance distribution show little to no correlation with halo mass: namely, clusters within the same halo mass bins can have significantly different amounts of Fe in their centres. We argue that this is related to their assembly and SMBH feedback histories, and we will expand on this in future works. Indeed, with the TNG model, simulated clusters have been shown to exhibit different central metallicity depending on their cool-core status (Vogelsberger et al., 2018), akin to what is suggested by observations. Finally, similar to the tails of cooler gas found in the temperature distributions, a lot of systems tend to have extended tails towards super-solar Fe abundances from gas cells that appear to be associated with the ISM or CGM of the BCG.
Overall, given the broadness of the Fe abundances distributions predicted by TNG-Cluster, the single-metallicity models typically assumed for fitting X-ray spectra are a poor approximation of the metal-enrichment complexity of the centres of clusters.
3.2 Inferring the ICM temperature distribution
| Model | Accepted fits | Acceptance rate (per cent) |
| bvapec | 322 | 91 |
| double bvapec | 154 | 44 |
| bvgadem | 160 | 45 |
| bvlognorm | 199 | 57 |
As shown and discussed above, as well as in Rohr et al. (2024) and Staffehl et al. (2025), the intrinsic temperature distribution of the ICM, as predicted by TNG-Cluster, is multiphase in nature. Therefore, an important aspect for the accuracy of spectral-emission models is their ability to reflect such a multiphaseness in their best fits – we tackle this next.
First, before quantifying the inference errors, we comment on how frequently it is even possible to successfully fit the X-ray spectra of XRISM/Resolve-like observations of the centres of clusters simulated in TNG-Cluster as a function of underlying spectral-modelling choices. Throughout, we adopt the mocking procedure described in Section 2.2 and the fitting analysis and operational definitions of “accepted” and “rejected” fits of Section 2.3.
Based on the success rates of the various models reported in Table 1, the bvlognorm model is the best performer among the multi-temperature ones, followed by the bvgadem model and the double bvapec, in that order. The bvapec model has, by far, the largest number of accepted fits: however, by assuming a delta function for the temperature distribution, it fails a priori to provide a realistic depiction of the ICM’s thermal properties. For the other models, the success rate can be taken as a measure for the compatibility between the intrinsic and the assumed temperature distributions. For example, the bvlognorm model assumes a more accurate depiction of the ICM temperature distribution as a unimodal distribution and remains insensitive to the sub-keV gas, which should have a negligible contribution in a XRISM/Resolve like spectrum. On the other hand, the bvgadem model, which tends to overestimate the contribution of the sub-keV gas, underperforms due to inflated distribution widths. With a similar acceptance rate, the double bvapec model also underperforms since the simulated clusters lack any visible bimodality, i.e. two similarly-prominent peaks.
Regarding accuracy, we assess first the best-fit values for the primary temperature value (or ICM effective temperature, Figure 5) and then, when appropriate, the width of the distribution (Figure 6). In both cases, we compare the best-fit parameters with the “true” values from the simulations by referring to both the mass-weighted gas properties of the ICM in a central sphere as well as the emission-weighted properties in a central pointing with XRISM-like FoV.
ICM effective temperature. For the “true” ICM single-value temperature we take the mean of the temperature distribution from the simulated clusters, i.e. the mean of each input distribution. This is compared against the best-fit values. The mode or median could also be good summary statistics of the TNG-Cluster temperature distributions, but they do not represent the outputs of the bvgadem and bvlognorm models. For the double bvapec model, we consider exclusively the best-fit primary temperature, namely the one that maximizes the relative contribution to the spectrum. We note that for a handful of systems, the double bvapec model finds a hotter, instead of a colder, secondary i.e. subdominant component temperature.
Based on Figure 5, top panels, all models tend to under-predict the mass-weighted mean temperature of the X-ray emitting gas that is actually located in the centre of each cluster. Whereas the bias is systematic, it is in fact relatively small. All models, on average i.e. across the TNG-Cluster sample, underestimate the temperature by 1.19 keV (an error of 17 per cent), but with a strong dependence on cluster mass. Namely, the best-fit temperature can underestimate the true one by about 1.5 keV (20 per cent) at the high-mass end, while being a reasonable approximation for lower-mass clusters. This result underscores the difficulty of assessing the true physical state of central ICM gas. In turn, assuming that TNG-Cluster returns reasonably realistic systems, the top panels of Figure 5 can provide guidelines to correct for this observational-analysis bias. This is the case on average, whereas for individual clusters the inference error can be much larger – we expand on outlier cases later.
In terms of emission-weighted quantities and comparing to the true properties of all the gas along the line of sight of a XRISM-like central pointing (Figure 5, bottom panels), the situation improves. All models produce a best-fit temperature that is indicative of the emission-weighted mean temperature of the gas for a large majority of the studied clusters, independent of halo mass. For the double bvapec model, in particular, and for both the top and bottom panels, the inference errors span the largest ranges, with a sizeable fraction of clusters whose observational-based (primary) temperature is substantially higher or lower than the true one. Moreover, the errors on the best-fit values for the double bvapec model are on average much larger than in any other case.
We find no significant differences in the quality of the inferences between the accepted and rejected fits, at least according to our definitions. In all cases, they occupy the same parameter space, suggesting that the rejection of the fit does not depend on the inference of the mean mass- or emission-weighted temperature.
ICM temperature width. Beyond the mean temperature, a key parameter in the description of the ICM temperature structure is the width of the distribution. As the “true” value we take the standard deviations of the best-fit Normal and Log-Normal distributions to the intrinsic gas (Figure 3 and Figure 4). We distinguish between the results for the mass-weighted gas in the 3D cluster centre versus emission-weighted gas for a central pointing with a XRISM-like FoV. We then compare these values to the inferred widths from the spectral fitting assuming, respectively, the bvgadem and bvlognorm spectral-emission models.
In order to assess both the accuracy and the significance of the observational inferences on the underlying thermal properties of the gas, in Figure 6 we show the residuals of the inferred thermal width as a function of the ground truth:
where and refer to the distribution width and its 1 error as inferred by fitting the spectra and where is the true width of the temperature distribution, as predicted by TNG-Cluster. These residuals so defined provide a gauge for the accuracy of the inferred temperature distribution width, at least for the accepted fits; on the other hand, the values in the rejected cases are primarily driven by the denominator and are representative of the higher uncertainties associated with a spectral fit that has not converged.
In Figure 6, we quantify the residuals for both the mass-weighted temperature distribution of the central gas (top) and the emission-weighted gas in a XRISM-like FoV (bottom). For both cases, the bvgadem model tends to overestimate the width of the distribution, with an average of 1.56 (-1.77,+0.55) and 1.27 (-1.67,+0.56), for top and bottom cases, respectively. On the other hand, the bvlognorm model tends to overestimate the average for the central mass-weighted gas (1.05 (-1.28,+0.70)), but is in a good agreement with the width of the emission-weighted gas in the XRISM FoV by an average of 0.41 (-1.05,+0.85). We note that, for both models, we find inferences that deviate by more than 5 from the true width of the distribution, even for accepted fits. These cases appear to be systems with temperature distributions that have pronounced tails of cooler or super-virial gas.
Overall, the bvlognorm spectral-emission model can better capture the true width of the broad temperature distribution of the central ICM gas, whether mass-weighted or emission-weighted and including projected contributions.
3.3 Inferring the ICM Fe abundance
We proceed to assess how well observational inferences of the ICM Fe abundance return the “true” underlying values, by following the same procedure as for the temperature, in fact by inspecting the results of the same fits of XRISM/Resolve-like spectra from TNG-Cluster discussed thus far.
In Figure 7 we show the comparison between the inferred Fe abundance from all spectral-emission models compared to the mass-weighted mean Fe abundance of the ICM in the actual 3D cluster centre (top) and the emission-weighted mean Fe abundance of gas from a central pointing with a XRISM-like FoV (bottom). As the top panel of Figure 7 shows, the inference on the Fe abundance is systematically different of the actual amount of Fe in cluster centres. Namely, all models underestimate the central Fe abundance by about 0.12 Solar (22 per cent). The offset varies with halo mass, being larger towards the low-mass end of our mass range. It is important to point out that such a systematic offset is significantly larger than the statistical errors on the Fe abundance typically quoted in the literature, of the order of 0.01, e.g. for the inferences from XRISM/Resolve observations available in the literature.
However, the inference error is at least partially driven by the subset of clusters that, despite being actually metal rich, are found to be significantly metal poorer when their X-ray spectra are analysed, irrespective of the adopted spectral-emission model. This discrepancy disappears if we account for emission (bottom panels), suggesting it is possibly driven by less X-ray bright gas. Interestingly, even though the rejected fits do not occupy a special region in the parameter space, as in the temperature case, a disproportional amount of clusters appear to belong to this subset.
Similar to the case of the emission-weighted mean temperature, all models consistently trace the emission-weighted Fe abundance, or at least they trace more accurately the emission-weighted than the mass-weighted Fe abundance, with no significant halo-mass dependence.
Overall, all models tend to underestimate the average Fe abundance of the gas in the centre of clusters. However, the single-point estimate for the amount of Fe is too simplistic of an approximation of the true, extended distribution. We speculate, for example, that extrapolating this value throughout the entire volume could still systematically inflate the total amount of Fe in the ICM.
4 Discussion
Overall, our results indicate that the assumptions on the temperature distribution of the gas may have an important impact on the inference of the properties of the ICM at the centre of clusters from X-ray spectral fitting, at least of those predicted by TNG-Cluster.
As shown in Figure 5 and Figure 7, the inferences of the bvapec model on the emission-weighted temperature and Fe abundance of the gas along the LoS are comparable to (if not more accurate on average than) the other spectral-emission models. We note that for the temperature, this is achieved despite us not using the so-called spectroscopic temperature weighting from Mazzotta et al. (2004), which was introduced to account for biases on the inference from CCD-resolution mock observations. Similarly, we find a good agreement between the true and inferred Fe abundance, even for haloes with temperatures between 2 and 3 keV, suggesting that our estimates are not subjected to the "inverse Fe-bias" (e.g. Rasia et al., 2008; Simionescu et al., 2009; Gastaldello et al., 2010). We believe that, in both cases, the bvapec model is able to provide a more accurate inference on the true temperature and Fe abundance of the ICM due to the higher spectral resolution, which reveals more lines that can place strong constraints on the model. However, our analysis remains limited to the cluster centres and the quality of the inference varies from target to target, suggesting that a broader examination of this topic, for a large sample of galaxy clusters over larger scales is necessary, but outside the scope of this project.
However, the bvapec model, by design, ignores the presence of multi-phase gas and cannot provide us any insights on the distribution of temperatures present in the ICM. Attempting to rectify this with the inclusion of another single-temperature component, in the form of the double bvapec model, does not improve the accuracy of the inferences on average (i.e. across the studied cluster sample) and actually returns a much larger fraction of clusters with very large inference errors on the ICM effective temperature, with biases in both directions.
The bvgadem and bvlognorm assume a Normal and Log-Normal distribution of temperatures, respectively, hence allowing a more complete description of the properties and diversity of the gas at the centre of clusters. For this reason, they return somewhat more frequently successful fits than the double bvapec model. And for the same reason, we argue that they should be prioritized, especially when attempting to study complex phenomena such as the effects of AGN feedback on the ICM, which could impact different gas phases differently (as is the case in the TNG model, see e.g. Truong et al., 2024). However, on average, the bvgadem and bvlognorm models do not necessarily return more accurate inferences of the ICM effective temperature and average Fe abundance. And, as shown, they too slightly albeit systematically underestimate the bulk properties of the ICM that is actually located close to the cluster centre.
In the following sections, we compare the fits produced by each model to the simulation-intrinsic distributions, in order to examine the quality of each inference in greater detail. We also investigate the impact that projection effects and brightness bias have on the inference of the properties of the central gas, in order to understand the systematic offsets between the two quantified so far. Finally, we also attempt to account for selection biases, such as focusing only on the X-ray brighter strong cool-core (SCC) clusters, in order to make our results more applicable to typical observational approaches.
4.1 Insights on the quality of the observational inferences, and of the fits, across spectral-emission models
As we have previously shown, all the spectral-emission models used in this work are capable of inferring with good accuracy the emission-weighted summary statistics of the cluster gas from a central pointing with a XRISM-like observation. Yet, for a large number of clusters, they fail to return an accepted best fit of the underlying X-ray spectrum. Moreover, as mentioned above and by comparing the filled and empty markers of Figure 5, Figure 6 and Figure 7, it would appear that the rejection of the X-ray spectra fits (at least based on our criteria, see Section 2.3.1) does not translate in much or overall worse inference errors. This can be appreciated via the quantitative summaries provided in Table 2 and Table 3. In the following, we hence provide further considerations as to why the quality of the fit may still be relevant when choosing which spectral-emission model to adopt to analyse the data.
4.1.1 Examples of true vs. inferred gas distributions
In Figure 8 we showcase the emission-weighted distributions of the ICM properties of a few example individual systems in comparison to their best-fit description, for all the spectral-emission models studied in this work. Each cluster is represented by two panels, one for temperature and one for Fe abundance, respectively, on top of each other. We show a total of 12 example clusters, selected based on the performance of the bvlognorm model to infer, respectively, effective temperature (top two rows), Fe abundance (bottom two rows; continued Figure) and temperature width (middle two rows).
We focus on the bvlognorm because it provides a good descriptor of the TNG-Cluster temperature distributions (subsection 3.1) and because it provides more frequently more accurate best-fit results for the temperature distribution widths (subsection 3.2). However, this discussion can be generalized for all spectral-emission models.
From left to right, we visualize the intrinsic and inferred gas property distributions for different example cases. Similar to Figure 3, we present the probability mass function of the emission-weighted gas distribution along the LoS as a fraction of the EM per bin. In the leftmost column, we show systems with the minimum absolute difference between the true and inferred value of the relevant property. We also showcase instances of an overestimation and underestimation, located at the 95 and 5 percentile of the sorted difference distribution, respectively (middle columns). Finally, we show examples of rejected fits (rightmost column), so that we can better see the intrinsic properties of the gas that, despite an accurate inference, led to model rejection.
Starting with the accepted spectral fits, we can understand how each model attempts to infer the input temperature distribution. The double bvapec model systematically infers temperatures that are not present or not associated with any minor peaks in the temperature distribution. Similarly, in many cases it fails to identify the most prominent peak, by inferring two temperatures around it. This suggests that the model, in an attempt to approximate a bimodal distribution of temperatures, does not provide an accurate representation of the actual distribution, but an attempt of residual minimisation.
The bvgadem model is capable of determining the peak of the temperature distribution, but systematically assumes a much larger amount of cooler gas compared to what is present. Finally, many cases show how the bvlognorm model provides a fairly accurate representation of the width of the temperature distributions. However, it tends to fail when the intrinsic temperature distribution deviates strongly from a Log-Normal distribution in the form of extended tails towards higher- or lower-temperature gas.
Regarding the Fe abundance distributions, all spectral-emission models make overly simplistic approximations. We see a range in the capability of all models to identify the mean Fe abundance, consistent with the trends found above.
Another important aspect regarding the inference of each model are the conditions that led to its rejection. As we can see in Figure 8 and as previously indicated in Figure 5 and Figure 7, even rejected fits of all models are capable of inferring a fairly accurate value of the mean temperature and Fe abundance emission-weighted distribution. For bvgadem and bvlognorm, the main and most frequent factor that leads to the rejection of the model is the width of the distribution: in some rejected cases, the inferred width is simply too large, infinity; in other rejected cases, the model infers a very narrow temperature distribution, i.e. with a poor constraint on the width itself.
4.1.2 Dependence of quality of the fit on number of photons
Considering the sensitivity of all models on the emission-weighted properties of the gas along the LoS, it is possible that a model cannot infer the temperatures of the less luminous gas contributing to the distribution because of a lack of counts.
In Figure 10 we examine this hypothesis by comparing each model inference on the temperature and Fe abundance as a function of the number of counts. It is evident from the distribution of our clusters that the rejected bvlognorm best-fits tend to have systematically lower counts than the accepted ones, but the count distributions of the two populations have a significant overlap. We note that the same behaviour holds for the accepted and rejected best-fits of the other spectral-emission models.
To understand the extent that exposure time may have on the rejection of our best-fits, we repeat the X-ray spectra mocking and analysis for the rejected clusters by increasing the exposure time to 300 ks, instead of the fiducial 100 ks adopted throughout. For the case of the bvlognrom model (which returns the largest success rate), we find that 55 per cent of the clusters that have been previously rejected turn out to have a valid fit with an increased exposure (not shown). This confirms the hypothesis that the rejection of a spectral fit for the fainter clusters in our sample is due to available exposure time.
However, this is not just a cluster-mass effect, as in fact there still remains a population of systems for which the best-fit is still rejected, even though their inferences on the emission-weighted mean temperature and Fe abundance are comparable with the rest of the sample. We find (although do not show) that clusters that are rejected even for a three-time longer exposure are either metal poor, with the majority of the emission-weighted mass in the ICM having an Fe abundance of 0.25 Solar, or deviate heavily from a Log-Normal distribution, with a significant amount of sub-keV gas. We note that such cases appear as almost separate populations, without any apparent correlation with halo mass (Figure 7, bottom left panel).
Regarding the systems that deviate from the Log-Normal distribution, it appears that they get rejected due to their poor goodness-to-fit. It is evident that the spectral-emission model’s parametrisation is not able to fully capture their temperature distribution, leading to poor fits, despite the increased number of counts. On the other hand, the rejection of the metal poor population is due to the fact that the spectral-emission models use the emission lines also to recover the temperature distribution of the gas. Whereas the shape of the continuum is a reliable diagnostic of the ICM’s bulk temperature, for plasma with temperatures higher than 2 keV, the emission-line ratios trace the presence of cooler or hotter gas, constraining the width of the distribution. However, if clusters have low elemental abundances, they will return low numbers of counts in spectral lines irrespective of the longer exposures, and in turn the modelling will be unable to provide any meaningful constraints on the width of the temperature distribution, leading to their rejection.
Overall this indicates that even the more accurate assumption of a temperature distribution is not a sufficient representation of the properties of the ICM, as predicted by TNG-Cluster, both for its thermal and more importantly its chemical structure. This highlights the limitations of current spectral-emission models and the need for more sophisticated descriptions of the intrinsic properties of the gas in order to better understand the ICM.
4.2 Effects of observational realism
Even though it might be tempting to assume that the inferred temperature and Fe abundance from a central-pointing spectrum are representative of the local conditions of the ICM, our analysis suggests otherwise. As shown e.g. in Figure 5 and Figure 7, we systematically find large inference errors on the mass-weighted state of the central ICM, but much smaller deviations from the emission-weighted properties of the projected gas along our LoS. Whereas this differentiated fit performance may not be surprising, it may have implications when searching for a physical interpretations of the data.
We now explore the individual sources of those differences and quantify their impact and systematic uncertainties. In the following sections we therefore attempt to understand and quantify the systematic uncertainties introduced by projection effects and the emission bias, by isolating the two.For the purposes of this analysis we focus again exclusively on the bvlognorm model as a reference point. However, the qualitative effects that we find are consistent across all models.
4.2.1 Projection effects
Due to inherent limitations of observations, it is not possible to differentiate emission based on the 3D position of its source. Therefore, a spectrum from a central pointing contains both the contribution from the centrally-located gas and of all the gas projected along the LoS. Previous works using simulated clusters have shown that projection effects can lead to severe misinterpretations of the intrinsic thermal and chemical properties of the ICM due to the gas along the LoS (e.g. Zhang et al., 2024; Štofanová et al., 2024).
In this section we attempt to quantify the impact of the projected gas in our inference errors by comparing the mass-weighted distributions of all the cluster gas in a central pointing with a XRISM-like FoV to the mass-weighted distribution of the gas cells that actually reside in the region of interest, in this case a sphere of a 33.5 kpc radius located in the centre of each cluster.
In Figure 11, we compare the temperature distribution of the gas in the centre of clusters to the total gas along the LoS observed in our mocks (top right), and their corresponding inference errors (top left). Since the temperature distribution of the gas in the clusters centre can be in most cases approximated by a unimodal distribution, we bin the clusters in percentiles based on the mode of their temperature distribution. In fact, due to the tight correlation between ICM temperature and the halo mass, the temperature bins of Figure 11 top right are comparable to mass bins. Differently than e.g. in Figure 5, here we isolate the effects of projections alone, by comparing directly mass-weighted vs. mass-weighted quantities rather than mass-weighted vs. emission weighted ones.
We see that the projection of gas from larger radii along the LoS shifts the temperature distribution towards lower temperatures (dashed vs. solid curves), due the projection of cooler gas from the outskirts of the cluster. The effect is more pronounced for systems with higher ICM temperature, or equivalently higher halo mass, due to the stronger temperature gradient up to in those systems. This is consistent with the strong dependence of the temperature inference on halo mass, found in Figure 5. By comparing the difference between the inferred temperature with the bvlognorm model on the one side and the the mean, mass-weighted temperature of the central vs. LoS gas, we find much larger inference errors in the former case, with average offsets of -1.21 keV (18 per cent) and 0.15 keV (3 per cent), respectively, with a strong dependence on halo mass. Given the significant reduction in inference errors between the two, we can conclude that projection effects dominate the observational bias.
Due to the projection of cooler gas, the width of the distribution systematically increases for all systems, but the effect shows a similar positive correlation with halo mass, being more pronounced for the most massive systems. The residuals on the inferred temperature width decrease by a factor of 4, from an average of 1.05 for the central gas to 0.26 for all the gas along the LoS. This suggests that the bvlognorm model is more representative of the distribution of temperatures along the LoS rather than the level of multi-phase gas in the central regions. However, given the inference errors, the statistical significance of the residuals effectively remains, on average, within 1 .
Projection effects are far more influential on the chemical structure of the ICM. In Figure 11 (bottom) we show the mass-weighted Fe abundance distributions of both central and projected gas. Similar to the temperature distributions, we bin the clusters based on the mode of the central gas, as it consists of one prominent peak. However, unlike the temperature of the ICM, TNG-Cluster shows no clear correlation between halo mass and metal enrichment. Our results indicate that, due to projection effects, a multitude of additional abundance peaks appear in the mass-weighted Fe abundance distribution of the gas across the LoS. These peaks consist of less metal-enriched gas and are more prominent in the systems with a higher core enhancement, since for the less enriched clusters the Fe abundance distribution remains effectively unimodal. The effect arises from large amounts of less-enriched gas that is located at larger galactocentric radii intersecting the LoS, which hence biases the mean Fe abundance low. Namely, projection effects change an average inference errors between the bvlognorm inferred metallicity and the intrinsic mass-weighted mean Fe abundance from -0.12 Solar (22 per cent) for the central gas to 0.08 Solar (27 per cent) for the full column. The effect has a strong dependence on the mean Fe abundance of the ICM, but it is independent of halo mass, explaining the trend we see in Figure 7.
Overall, our results indicate that systematic errors in the inferences of the temperature and metallicity of the ICM at the centre of clusters can be greatly reduced by de-projecting the observational results, irrespective of the adopted spectral-emission model. However, we note that such an approach would require observations and analysis of spectra from the full extent of the cluster, which can not be performed within the purview of a single pointing. Still, the full minimization of projection effects may still be hampered by the assumptions on e.g. the temperature distribution of the ICM along the LoS and the geometry of the cluster.
4.2.2 Effects of mass vs. emission weighting
Previous studies have shown that observed ICM values are more related to emission-weighted properties than to e.g. mass-weighted properties (e.g. Biffi et al., 2017, 2018). An emission spectrum, and models to fit emission spectra, are necessarily sensitive in a weighted way towards the most luminous gas, rather than the entire mass of the ICM. As a result, physical properties of the ICM inferred by X-ray spectroscopy, such as the amount of Fe in the ICM, can be biased by the properties of the brightest gas.
In this section, we attempt to quantify how the brightest gas along the LoS contributes to the inference errors on the properties of the gas by comparing the mass- and emission-weighted distributions of the gas along the LoS, and not just that of the central sphere, as well as how it influences inferences by the bvlognorm model.
Starting with the temperature, Figure 12 shows the difference between emission- and mass-weighted distributions (top right). Overall, these are small, and primarily visible in the sub-keV tails. This is expected as the associated gas is located close to the cluster centre, is densest, and is therefore brighter than most other gas along the line of sight. The influence on the quality of the bvlognorm inference is similarly mild (top left). We find a shift from an average inference error of 0.15 keV (3 per cent) to -0.08 keV (1 per cent) from the mass-weighted to the emission-weighted ICM effective temperature.
The effect on the the inference of the width of the temperature distribution due to this brightness bias is relatively small, with the residuals increasing to an average of 0.4, from 0.26 for the mass-weighted case. This further highlights the previously-found trend of the Log-Normal distribution remaining unaffected by the enhancement of the sub-keV gas. (not shown)
Finally, in Figure 12 (bottom) we examine the effects of the brightness bias on the Fe abundance distribution (bottom right). Similar to the temperatures, by favouring the brightest gas along the line of sight, there is an enhanced contribution from the super-Solar tails, that are associated with the cluster centre, and a relatively suppressed contribution from gas in the outskirts of the cluster. As a result, the emission-weighted mean Fe abundance is biased high, resulting in an average inference error of 0.08 Solar (27 per cent) compared to the average 0 Solar (<1 per cent), inferred by the bvlognorm model for the mass-weighted case.
Overall, for central pointings, the brightness bias has a minimal influence on the accuracy of the observational inferences on the properties of the projected gas. This could become more a significant issue for off-centre pointings, where the projected gas becomes comparable to the medium that is targeted. Examining and quantifying the influence of those biases in such cases is left for a future study.
4.3 Selection effects: i.e. the case of cool-core clusters
In our analysis, we have been considering the entire sample of 352 clusters in TNG-Cluster at . However, samples of observed clusters tend to not be as diverse, due to selection effects. For example, many of the clusters in the literature are cool core (CC), as they are significantly brighter than their non cool-core (NCC) counterparts and considered the best targets for the study of AGN feedback. Therefore, we now investigate if the inferences of our models depend on the cool-core state of a cluster.
In Figure 13 we present the offsets between inferred and mean, mass-weighted temperature and Fe abundance for the gas in the centres of our clusters, as a function of their central cooling time and central entropy (from Lehle et al., 2024). For all cases, we find no strong correlation between the inference error and the cool core metric, although the scatter is larger for weak cool core (WCC) and NCC clusters.
Regarding the rejected best-fits, while a number of strong CC (SCC) clusters fail to produce an accepted fit for all models, the rejected fits disproportionally arise from the NCC subsample. Namely, 70 per cent of the SCC subsample produces an accepted double bvapec and bvgadem best-fit and 80 per cent a valid bvapec and bvlognorm; these figures are much higher than the general population acceptance rates. We note that, accounting only for SCC, the bvlognorm model has the best acceptance rate, surpassing even the single temperature bvapec one. As previously shown, constraining the temperature distribution depends on the number of counts and therefore it can be more easily accomplished the brighter the associated gas is.
Finally, we examine the associated observational inferences on the properties of the ICM, if our sample of clusters consists only of SCC clusters, that are easier to detect, compared to their NCC counterparts, due to their apparent X-ray brightness (e.g. Eckert et al., 2011). Selecting only for the systems with central cooling times below 1 Gyr (following Lehle et al., 2024), we examine their impact on temperature and Fe abundance.
In Figure 14 we show the distribution of the CC clusters versus the full sample (shown in Figures 5 and 7, for comparison). For both the temperature and Fe abundance, the average inference error increases or remains similar to that of the entire sample, respectively. For example, for the accepted bvlognorm fits, the change is from -1.21 keV (18 per cent) to -0.68 keV (14 per cent) and from -0.12 Solar (22 per cent) to -0.05 Solar (9 per cent) for the mean, mass-weighted temperature and Fe abundance in the central sphere, respectively. However, excluding the faintest objects does remove the majority of the outliers, suggesting that the least accurate inferences occur for NCC clusters.
5 Summary and conclusions
In this paper, we have used the outcome of TNG-Cluster, a cosmological magnetohydrodynamical simulation of 352 massive clusters and their galaxies, to assess the accuracy with which it will be possible to infer the temperature and metallicity of clusters’ centres via high-resolution X-ray spectroscopy.
To this end, we have quantified the thermal and chemical properties of the ICM at the centre of TNG-Cluster systems at and we have developed an end-to-end pipeline to generate and analyse synthetic X-ray spectra that could be obtained with XRISM/Resolve-like observations targeting the core regions of clusters at . We have hence quantified the ability of commonly-used spectral-emission models (bvapec, double bvapec, bvgadem, and bvlognorm) to recover the intrinsic temperature distribution and Fe abundance of the gas.
Our main results are as follows:
The intrinsic temperature distribution of the gas in the centre of clusters, both mass-weighted and emission-weighted, can not be completely captured by either a Normal or a Log-Normal distribution. More complicated functional forms may be necessary to accurately describe the temperature distribution of the ICM in the non-negligible fraction of systems that exhibit important tails towards both cooler and super-virial temperatures (Section 3.1, Figure 3 and Figure 4).
The central ICM temperature inferred via all spectral-emission models is consistent with the true emission-weighted average along the line of sight, but systematically underestimates the mass-weighted temperature of the gas that is truly located in the clusters centre. This inference bias reads, on average (i.e. across the whole population of 352 clusters), -1.2 keV (17 per cent; -0.73,+0.72 keV for the 16th-84th percentiles), so it is arguably small, but has a dependence on halo mass (Section 3.2, Figure 5). For clusters, all spectral-emission models return mean temperatures that are smaller by 1.5 keV, on average, than the true mass-weighted average one.
The width of the temperature distribution inferred by the bvlognorm model, is, on average, overestimated compared to the true mass-weighted distribution of the central ICM, but within 1 of the emission-weighted distributions. In contrast, the bvgadem model overestimates the width of both mass- and emission-weighted temperature distributions to a larger extent (Section 3.2, Figure 6).
The inferred Fe abundance is consistent with the true mean, emission-weighted Fe abundance of the gas along the line of sight, i.e. for a central pointing with a XRISM-like FoV. However, all models similarly underestimate the mean Fe abundance in the cluster centre by about 0.12 Solar (22 per cent), which is significantly larger than the statistical uncertainties usually associated to the measurements (Section 3.3, Figure 7).
The rejection of fits of the X-ray spectra correlates with the number of collected photon counts but, interestingly, not with errors in the inference. More than half of the initially rejected bvlognorm fits do succeed with longer exposure times. However, metal-poor systems and systems with temperature distributions that deviate strongly from a Log-Normal, fail to produce an accepted spectral fit, even if we account for longer exposures (300ks vs 100ks; Section 4.1, Figure 10).
When considering only strong cool-core clusters, as is the case in many observational samples, model inferences on central ICM temperature improve only marginally for the average cluster (with e.g. inference errors of 0.7 keV on average compared to 1.2 keV across a randomly-selected overall sample). The CC-selection effect on the inference of Fe abundance is even smaller. The main advantage of selecting for SCCs is the removal of strong outliers, i.e. of clusters for which the observational inferences of ICM properties are catastrophically wrong (Section 4.3, Figure 14).
Overall, our analysis highlights the importance of properly modelling the complex temperature and metallicity distributions of the ICM at the centre of clusters. The different spectral-emission models that we have considered have different strengths and weaknesses, or practical justifications. Barring the double bvapec, which returns highly inaccurate results for more clusters than the other models, all studied assumptions return overall similarly accurate inferences of the effective temperature and average metallicity of the ICM at the centre of clusters. However, based on the predictions of TNG-Cluster, we recommend to use models that allow for a temperature distribution, as both single and double temperature models simply fail at capturing the diversity and complexity of the ICM. Similarly, the single-value Fe-abundance assumption, typical of all ICM X-ray spectra analyses, seems utterly inadequate to represent the multifacetedness of the ICM enrichment, even in the innermost regions of clusters – a qualitative shift in the way ICM X-ray spectra are fitted for, and hence interpreted, seems of the essence. Finally, due to the large influence of the effects associated with observational realism, particularly projection effects, a careful deprojection to the 3D properties of the gas from observed data is required.
While we have primarily focused on the ICM Fe abundance, in future work we will assess the accuracy of XRISM/Resolve-like observations and analyses in recovering the abundances of other metal species, in addition to gas kinematics. It remains to be quantified how the figures put forward here translate to lower, CCD-like resolution spectra, such as with XMM-Newton, to observations targeting off-centre pointings and cluster outskirts, and to science cases that rely on spatially-resolved analyses.
Data Availability
All the output of the TNG-Cluster simulation is now publicly available, as are the IllustrisTNG simulations: see www.tng-project.org/data and Nelson et al. (2019). The data generated in this project, including XRISM Resolve mocks, spectra, and spectral-fitting results, will be made available upon acceptance.
Acknowledgements
We thank the reviewer for their constructive and useful suggestions. DC and AP thank Joey Braspenning for useful comments on the paper. DC is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD) and, together with AP, acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich). DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1).
The TNG-Cluster simulation suite has been executed on several machines: with compute time awarded under the TNG-Cluster project on the HoreKa supercomputer, funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research; the bwForCluster Helix supercomputer, supported by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG; the Vera cluster of the Max Planck Institute for Astronomy (MPIA), as well as the Cobra and Raven clusters, all three operated by the Max Planck Computational Data Facility (MPCDF); and the BinAC cluster, supported by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG. The material is based upon work supported by NASA under award number 80GSFC24M0006.
All the analysis and computations associated to this paper have been realized on the Vera cluster of the MPCDF.
Appendix A spectral-emission model systematic uncertainties
This section reports in detail, in a series of Tables (2, 3), the systematic uncertainties, associated with the choice of spectral-emission model, on the inference of the true properties of the ICM, as predicted by TNG-Cluster.
| Inferred - True value | bvapec | double bvapec | bvgadem | bvlognorm |
| Effective Temperature (kT) [keV] | ||||
| Mass-weighted central gas (all) | -1.05 (-0.75,+0.74) | -1.32 (-1.07,+1.22) | -1.18 (-0.78,+0.74) | -1.21 (-0.73,+0.72) |
| Mass-weighted central gas (accepted only) | -1.09 (-0.72,+0.72) | -0.68 (-0.92,+0.87) | -0.98 (-0.59,+0.54) | -1.17 (-0.67,+0.68) |
| emission-weighted FoV gas (all) | 0.09 (-0.24,+0.16) | -0.17 (-0.58,+0.86) | -0.05 (-0.17,+0.22) | -0.08 (-0.22,+0.24) |
| emission-weighted FoV (accepted only) | 0.06 (-0.21,+0.15) | 0.22 (-0.95,+0.69) | 0.00 (-0.16,+0.11) | -0.15 (-0.17,+0.14) |
| Temperature width Residuals () (2) | ||||
| Mass-weighted central gas (all) | 1.56 (-1.77,+0.55) | 1.05 (-1.28,+0.70) | ||
| Mass-weighted central gas (accepted only) | 2.08 (-1.73,+0.64) | 1.51 (-1.21,+0.79) | ||
| emission-weighted FoV gas (all) | 1.27 (-1.67,+0.56) | 0.41 (-1.05,+0.85) | ||
| emission-weighted FoV gas (accepted only) | 1.83 (-1.58,+0.81) | 0.81 (-0.89,+0.92) | ||
| Fe abundance [Solar] | ||||
| Mass-weighted central gas (all) | -0.13 (-0.11,+0.10) | -0.12 (-0.11,+0.09) | -0.12 (-0.12,+0.09) | -0.12 (-0.12,+0.09) |
| Mass-weighted central gas (accepted only) | -0.14 (-0.11,+0.10) | -0.09 (-0.05,+0.07) | -0.08 (-0.04,+0.06) | -0.08 (-0.04,+0.07) |
| emission-weighted FoV gas (all) | -0.02 (-0.02,+0.03) | -0.01 (-0.02,+0.03) | -0.01 (-0.02,+0.03) | -0.00 (-0.02,+0.03) |
| emission-weighted FoV gas (accepted only) | -0.02 (-0.02,+0.03) | 0.01 (-0.01,+0.03) | 0.01 (-0.01,+0.03) | -0.01 (-0.01,+0.03) |
| Inferred / True value | bvapec | double bvapec | bvgadem | bvlognorm |
| Effective Temperature (kT) [keV] | ||||
| Mass-weighted central gas (all) | -0.14 (-0.08,+0.06) | -0.19 (-0.12,+0.17) | -0.18 (-0.06,+0.09) | -0.18 (-0.08,+0.07) |
| Mass-weighted central gas (accepted only) | -0.15 (-0.07,+0.06) | -0.11 (-0.18,+0.16) | -0.15 (-0.06,+0.06) | -0.19 (-0.08,+0.08) |
| emission-weighted FoV gas (all) | 0.05 (-0.07,+0.02) | -0.02 (-0.14,+0.20) | -0.01 (-0.03,+0.04) | -0.01 (-0.05,+0.04) |
| emission-weighted FoV gas (accepted only) | 0.04 (-0.06,+0.01) | 0.06 (-0.23,+0.17) | 0.00 (-0.03,+0.02) | -0.03 (-0.04,+0.03) |
| Fe abundance [Solar] | ||||
| Mass-weighted central gas (all) | -0.25 (-0.19,+0.17) | -0.22 (-0.20,+0.20) | -0.23 (-0.20,+0.20) | -0.22 (-0.20,+0.17) |
| Mass-weighted central gas (accepted only) | -0.26 (-0.19,+0.16) | -0.16 (-0.13,+0.13) | -0.15 (-0.11,+0.10) | -0.15 (-0.12,+0.11) |
| emission-weighted FoV gas (all) | -0.04 (-0.08,+0.04) | 0.00 (-0.09,+0.09) | -0.01 (-0.09,+0.09) | 0.01 (-0.08,+0.10) |
| emission-weighted FoV gas (accepted only) | -0.03 (-0.07,+0.07) | 0.03 (-0.05,+0.07) | 0.03 (-0.04,+0.06) | 0.05 (-0.06,+0.07) |
Appendix B Inferred LoS velocity dispersion
In this section, we report, in addition to the temperature and metallicity properties of the ICM – the focus of the paper –, also the LoS velocity dispersion values inferred according to each spectral-emission model. We do so for the full sample of studied clusters. As previously mentioned, comparing these values to the intrinsic 3D velocity dispersion of the gas cells from the simulation exceeds the scope of this work. However, considering potential biases due to degeneracies among all free parameters and considering the current interest in the kinematics of cluster centres enables by XRISM, we want to provide results also on our kinematic inferences.
In Figure 15 we present the inferred velocity dispersions for all spectral-emission models used in this work as a function of redshift. We note that we find little to no difference in the inferred values between models, with the possible exception of the double bvapec model for low mass halos. This suggests that the inferred kinematics appear to be, for the most part or at least for the average cluster, unaffected by the parametrisation of the ICM’s thermal distribution.
For comparison, we also include in Figure 15 the observationally-inferred velocity dispersions from the cores of a few selected galaxy clusters, as available from the current literature (large black markers, from lower to higher masses: Centaurus, Hydra A, Perseus, Abell 2029, Coma and Ophiuchus). We see that observations are consistent with the TNG-Cluster predictions, deviating by less than 1 from the average trends. The only outlier appears to be Ophiuchus, at the highest mass end. However, in the observational setup, the central region only covers the central 25 kpc, compared to the 33.5 kpc used in this work, and the velocity dispersion appears to increase with radius (Fujita et al., 2025). This might imply that the velocity dispersion of Ophiuchus may also fall within 1 of the TNG-Cluster predictions, if the central pointings were matched. This preliminary and at-face-value comparison may serve as a useful starting point for future, more tailored analyses (e.g. Xrism Collaboration et al., 2025b, c).
Appendix C On fitting the temperature distributions of simulated ICM
As discussed in Section 3.1, we fit the intrinsic temperature distribution of the ICM predicted by TNG-Cluster with both a Normal and a Log-Normal distribution. In Figure 16 we present the residuals, i.e. the difference between intrinsic and modelled probability mass functions, for both fit results. On average (thick curves), the offsets do not exceed the 1 per-cent level, suggesting that both Normal and a Log-Normal functional forms can provide a reasonable description of the thermal properties of the gas across our entire cluster sample.
However, the tensions are more significant for individual halos (thin curves), with a large cluster-to-cluster variation reflected in the large scatter of values. As we have already highlighted for the entire sample, the tensions primarily arise from highly asymmetric intrinsic distributions, due to extended tails of either hotter or cooler gas, relative to the mode of the distribution. We note that for both functional forms, we find an increase in the residuals on the sides of the temperature peak due to over-/under-fitting, as explained in the text.